x + -3(x + 3) >= 77
Answer by lenny460(1073)

x + -3(x + 3) > = 77

-3 * x = -3x
-3 * 3 = -9
Therefore:
x + -3x - 9 > 77
Combine like terms:
x + -3x = -2x
-2x - 9 > 77
-2x - 9 + 9 > 77 + 9
-2x > 86

-2x/-2 > 86/-2
Answer:
x < -43




Lennox Obuong
Algebra Tutor
Nairobi, Kenya